# Smart Expert System: Large Language Models as Text Classifiers

### v1

@@ VI-B Fine-tuning strategy
More importantly, after fine-tuning, the U/E value cross models and datasets are down to 0, except for the Llama-3-8B(F) in tweet classification. This improvement in the standardized output makes the result consistent and makes the smart expert system more reliable.
@@ VI-C Limitations
While LLMs demonstrated impressive proficiency in text classification across various domains and sentiment analysis, our experiments also uncovered a range of limitations that affect their practicality as zero-shot text classifiers. These constraints became particularly evident when juxtaposing LLMs with traditional ML algorithms and NN architectures.
  • Inconsistency in Standardized Output Formats: One key limitation of LLMs is their inconsistent output format. This can hinder the seamless integration of LLMs into existing systems that rely on specific output formats. For instance, when expecting results in a standardized JSON format, the LLMs may occasionally deviate by providing unstructured text strings or introducing variations in the keys or values within the JSON output. This inconsistency poses a challenge for downstream applications that require consistent and structured data for further processing.
  • Restrictions on Content Classification: Another limitation is the potential reluctance of certain LLMs to classify specific types of content. This reluctance can stem from various reasons, such as containing sensitive or impolite content. Furthermore, some LLMs may decline to classify content that exceeds their processing capabilities. Such limitations can constrain the scope of applications where LLMs can be effectively deployed, particularly in diverse or nuanced textual content scenarios.
  • Proprietary Model Constraints: Commercial closed-source LLMs introduce several operational constraints that can impede scalability and accessibility. API request rate limits service providers impose can throttle throughput and affect real-time application performance. Additionally, the cost associated with API calls can become prohibitive at scale, especially when dealing with large volumes of data requiring classification.
  • Hardware Resource Demands: The computational demands of utilizing LLMs, especially larger models, can impose significant hardware requirements. This includes the need for robust CPU and GPU resources to support the intensive processing capabilities of LLMs. The scalability and deployment of LLMs may be hindered by these hardware constraints, particularly in environments where access to high-performance computing resources is limited or costly. Detailed insights into the specific hardware prerequisites can offer valuable guidance for practitioners seeking to leverage LLMs effectively.
  • Time-intensive Processing: LLMs often exhibit longer inference times than traditional ML or NN models. The complex architecture and extensive parameterization of LLMs contribute to the prolonged processing duration required for generating classification results. This time-consuming nature can impact real-time or high-throughput applications requiring rapid responses, highlighting the trade-offs between accuracy and efficiency when employing LLMs for text classification tasks.
@@ VII Conclusion and Future Work
In conclusion, our study has demonstrated the potential of LLMs as effective text classifiers within Smart Expert Systems, often surpassing traditional ML and NN approaches. Despite some LLMs showing lower accuracy in certain cases, strategic fine-tuning has proven to be a powerful method to enhance their domain-specific performance.
Our study underscores the transformative impact of LLMs on text classification, particularly in their ability to streamline the process by obviating the need for labor-intensive data preprocessing and bespoke model development. This characteristic is especially beneficial for small businesses seeking cost-effective solutions to integrate intelligent text classification into their operations without requiring deep expertise in ML or DL. The implications of our findings are profound: LLMs not only democratize access to cutting-edge AI technology but also empower organizations with limited resources to leverage sophisticated NLP tools. By harnessing the power of LLMs, businesses can efficiently process user feedback, enhance spam detection mechanisms, and automate workflows with minimal engineering effort.
For future work, we aim to address the challenges associated with fine-tuning, such as the need for substantial computational resources and expertise. Alternative strategies to improve LLM performance could include more detailed background provisions and precise label definitions. Refining non-standardized outputs remains a priority; employing a secondary LLM to process initial results could reduce user effort rates and improve overall system efficiency. As we continue refining these models, we aim to make LLM-powered expert systems more accessible and user-friendly across various applications.

### v2

@@ VI-B Fine-tuning strategy
More importantly, after fine-tuning, the U/E value cross models and datasets are down to 0, except for the Llama-3-8B(F) in tweet classification. This improvement in the standardized output makes the result consistent and makes the system more reliable.
@@ VI-C Limitations
While LLMs demonstrated impressive proficiency in text classification, our experiments also uncovered a range of limitations when leverage LLMs as text classifiers.
  • Inconsistent Output Formats: LLMs often produce inconsistent output formats, which can disrupt the integration into systems requiring standardized results (e.g., JSON format). This inconsistency challenges downstream applications that depend on structured data.
  • Content Classification Restrictions: Some LLMs may refuse to classify certain types of content due to sensitivity or processing limitations, restricting their application scope in diverse or nuanced scenarios.
  • Proprietary Model Constraints: Closed-source LLMs can limit scalability due to API rate limits and potentially prohibitive costs associated with high-volume usage, affecting real-time performance and accessibility.
  • Hardware Demands: Utilizing LLMs, particularly larger models, requires significant CPU and GPU resources. This can hinder scalability and deployment in environments with limited access to high-performance computing.
  • Time-intensive Processing: LLMs typically have longer inference times, impacting real-time or high-throughput applications. This trade-off between accuracy and efficiency is crucial for practitioners to consider.
@@ VII Conclusion and Future Work
In conclusion, our study has demonstrated the potential of LLMs as effective text classifiers, often surpassing traditional ML and NN approaches. Strategic fine-tuning has proven to be an influential method for enhancing LLMs’ domain-specific performance.
Our findings highlight the adaptability of LLMs in streamlining the text classification process by eliminating the need for extensive data preprocessing. This adaptability is particularly beneficial for small businesses looking for cost-effective solutions to integrate intelligent text classification without the requirement for deep ML or DL expertise. By democratizing access to advanced AI technology, LLMs empower organizations with limited resources to leverage sophisticated NLP tools. Businesses can efficiently process user feedback, enhance spam detection mechanisms, and automate workflows with minimal engineering effort, demonstrating the reliable and high-performance standards of our approach.
For future work, we aim to focus on making the system more reliable. Directions include but are not limited to employing a secondary LLM to process initial classification results, which could reduce U/E rates.
Acknowledgements
This study is supported by the U.S. National Science Foundation under grant Nos. IIS-2236579, IIS-2302786 and IOS-2430224.